Week of 190506

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • local: Miro (Chair, DB), Maarten (ALICE), Julia (WLCG), Belinda (Storage), Gavin (Compute), Alberto (Monitoring), Ivan (Atlas), Vladimir (LHCb)
  • remote: Alexander (NL-T1), Dave (FNAL), Jens (NDGF), Xin (BNL), Marcelo (CNAF), Xavier (KIT), Di (TRIUMF), Victor (JINR), Christoph (CMS)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity
      • User jobs, MC productions, and staging this week
    • Issues
      • CERN:
      • RAL:
        • Continuing migration from Castor to ECHO

Sites / Services round table:

  • ASGC: NC
  • BNL: done with transition of job environment to SL7 on the ATLAS T1 computing farm, transparent to users.
  • CNAF: NTR
  • EGI: NC
  • FNAL: Downtime on Tues to enable dual stack on both dCache instances
  • IN2P3: IN2P3-CC will be in maintenance on June 4th, a Tuesday. As usual details will be available one week before the event.
  • JINR: We offer 10000 core/1000 10-core slots in the farm for CMS from now. LHCOPN has been upgraded to 100Gbps.

  • KISTI: NC
  • KIT: NTR
  • NDGF: No issues to report. Planning for a dCache 5.1 upgrade next week unless problems are found on the testing system.
  • NL-T1: Last Wednesday there was a short network maintenance at SURFsara that went well and was completely transparent but unfortunately due to an internal miscommunication we were only able to announce it about 14 hours in advance.
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: NC
  • TRIUMF: Scheduled power outage at the building where the old data center is located and the old computing cluster is in downtime for about one week, we will migrate them to SL7 as bare-metal WN after coming back.

  • CERN computing services: NTR
  • CERN storage services: NTR
  • CERN databases: NTR
  • GGUS: NTR
  • Monitoring:
    • Draft availability/reliability reports for April sent around
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR

AOB:

Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2019-05-20 - DanieleCesini
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback